### v1

# Enhancing deep neural networks through complex-valued representations and Kuramoto synchronization dynamics

Abstract
Neural synchrony is hypothesized to play a crucial role in how the brain organizes visual scenes into structured representations, enabling the robust encoding of multiple objects within a scene. However, current deep learning models often struggle with object binding, limiting their ability to represent multiple objects effectively. Inspired by neuroscience, we investigate whether synchrony-based mechanisms can enhance object encoding in artificial models trained for visual categorization. Specifically, we combine complex-valued representations with Kuramoto dynamics to promote phase alignment, facilitating the grouping of features belonging to the same object. We evaluate two architectures employing synchrony: a feedforward model and a recurrent model with feedback connections to refine phase synchronization using top-down information. Both models outperform their real-valued counterparts and complex-valued models without Kuramoto synchronization on tasks involving multi-object images, such as overlapping handwritten digits, noisy inputs, and out-of-distribution transformations. Our findings highlight the potential of synchrony-driven mechanisms to enhance deep learning models, improving their performance, robustness, and generalization in complex visual categorization tasks.
1 Introduction
Learning structured representations in artificial neural networks (ANNs) has been a topic of extensive research (Zhang et al., 2013; Chiou, 2022; Dittadi, 2023; Le Khac, 2024), yet it remains an open challenge (Schott et al., 2021; Dittadi, 2023). Notably, some researchers argue that the inability of ANNs to effectively bind and maintain structured representations may underlie their limited generalization capabilities and susceptibility to distributional shifts (Greff et al., 2020).
In neuroscience, the Binding Problem (Treisman, 1996; Roskies, 1999; Singer, 2007) refers to the brain’s capacity to integrate various attributes of a stimulus—such as color, shape, motion, and location—into a unified perception. This process involves understanding how distinct features of an object are combined across different processing stages, enabling the brain to construct meaningful and cohesive representations of individual objects within a scene. Neural synchrony has been proposed as a key mechanism underlying this integrative process (Singer, 2007; Uhlhaas et al., 2009).
Kuramoto dynamics and related oscillator models have been widely employed in computational neuroscience to explore synchronization phenomena in neural systems (Breakspear et al., 2010; Chauhan et al., 2022). These models provide insights into complex neural processes, such as phase synchronization and coordinated neural activity. Beyond neuroscience, the Kuramoto model has also found applications in artificial intelligence, offering a framework for understanding synchronization in complex systems (Ódor & Kelling, 2019; Rodrigues et al., 2016). Recently, its utility has extended to computer vision tasks (Ricci et al., 2021; Miyato et al., 2024), demonstrating its potential to enhance representation learning in ANNs.
Building on these insights, we propose leveraging the Kuramoto model to investigate the role of neural synchrony in convolutional neural networks (CNNs) for multi-object classification. We hypothesize that incorporating neural synchrony, inspired by the brain’s solution to the Binding Problem, can be implemented using Kuramoto dynamics within ANNs, thereby enhancing their generalization abilities.
To test this hypothesis, we design a hierarchical model, KomplexNet, that integrates layers of complex-valued units with a bottom-up information flow. In KomplexNet, Kuramoto dynamics are applied at the initial layer to induce a synchronized state, which is then propagated through subsequent layers via carefully designed complex-valued operations. This approach enables the model to exploit the phase dimension of its neurons to bind visual features and organize visual scenes into distinct object representations, while the amplitude dimension retains standard CNN functionality.
We further extend KomplexNet by incorporating feedback connections to refine synchronization through top-down information. This extension demonstrates the critical role of top-down processes in enhancing phase synchrony and structuring object representations. Overall, our findings highlight the potential of neural synchrony mechanisms, modeled using Kuramoto dynamics, to improve the robustness, generalization, and representational capacity of deep learning architectures.
Overall, our contributions are as follows:
We introduce KomplexNet, a complex-valued neural network that leverages Kuramoto dynamics for multi-object classification.
KomplexNet has better classification accuracy than comparable baselines.
KomplexNet also exhibits better robustness to images perturbed with Gaussian noise and generalization to out-of-distribution classification problems.
Extending KomplexNet with feedback connections leads to better phase synchrony, exhibiting higher robustness and generalization abilities than KomplexNet without feedback.
2 Related work
Complex-valued models.
Complex-valued neural networks are popular and extensively used in signal processing to model complex-valued data, such as spectrograms (see (Bassey et al., 2021) for a review). The term Complex-Valued Neural Network (CVNN) is commonly used to refer to fully complex networks: not only is the activation function complex, but so are the parameters. Trabelsi et al. (2017) proposes a list of operations adapted to a parametrization in the complex domain, including convolutions, activation functions, and normalizations. Moenning & Manandhar (2018) systematically compares complex-valued networks and their real-valued counterparts for object classification. Their findings highlight the importance of the choice of activation functions and architectures reflecting the interaction of the real and imaginary parts. However, when applied to real-valued data, the field lacks appropriate conversion mechanisms to the complex domain. One proposal by Yadav & Jerripothula (2023) includes a novel way to convert a real input image into the complex domain and a loss acting on both the magnitude and the phase. They implement their transformation on several convolutional architectures and outperform their real-valued counterpart on visual categorization datasets. None of these papers use phase synchrony as a mechanism for perceptual organization in multi-object scenes.
Synchrony in artificial models.
Some work has explored binding by synchrony or leveraging synchrony in artificial models without complex-valued activity. Ricci et al. (2021) proposes a framework for learning in oscillatory systems, harnessing synchrony for generalization. This approach is, however, limited by its learning procedure: the model is designed to learn to segment one half of an image and generalize on the other. Zheng et al. (2022) extends a spiking neural network with attention mechanisms to solve the binding problem, though the impact of synchrony on performance or robustness is not evaluated. While these models successfully group entities by synchronizing the spikes of neurons, synchrony was not designed to assist in solving visual tasks. In particular, it is unclear from the work if and how the resulting representations help improve the neural network’s overall performance, robustness, or generalization ability.
Complex-valued representations and binding by synchrony.
A growing body of literature uses complex-valued representations to explicitly model neural synchrony. Early models were designed to implement a form of binding by synchrony via complex-valued units to perform phase-based image segmentation (Zemel et al., 1995; Weber & Wermter, 2005) or object-based attention (Behrmann et al., 1998). These models were shallow architectures, and they were trained on small datasets. Different mechanisms, including feedback mechanisms (Rao et al., 2008; Rao & Cecchi, 2010; 2011), were later explored to influence synchrony in deeper architectures. However, all models were trained on datasets that remained limited to toy objects. Specifically, despite the simplicity of the objects, the images contained individual objects only, limiting the potential benefit of synchrony. Finally, Reichert & Serre (2013) scaled to Boltzmann machines and multi-object datasets. The authors proposed a general framework that included operations for binding by synchrony. Binding was shown to emerge through the phase of neurons. However, the approach did not include backpropagation training or end-to-end deep learning. Specifically, a real-valued Boltzmann machine was first trained, and the phases were introduced during test. Synchrony was, therefore, a completely emergent property and did not take any part in helping the model learn the task. Building on this work, Löwe et al. (2022) adapted the approach for training a complex auto-encoder to reconstruct multi-object images. This model was fully complex, even during training. The phase synchrony helped the model reconstruct an input image and outperform its real-valued counterpart. Finally, Stanić et al. (2023) scaled up the model to more objects and color images by adding a contrastive objective on the phases, followed by Gopalakrishnan et al. (2024) who improved the phase synchrony using recurrence and complex-weights. Our work distinguishes itself from previous work, notably in how we exploit synchrony mechanisms. In all the aforementioned approaches, synchrony is sought as an emergent property of the task and the neural operations implemented in the network. In contrast, we propose to introduce it using a Kuramoto system as an explicit synchronizer. We then study the benefit of synchrony for object categorization performance, as well as robustness and generalization.
3 Binding by synchrony and Gestalt criteria
The binding by synchrony theory, as supported by a body of research both experimental (Singer, 2007; von der Malsburg, 1981) and computational Milner (1974); Grossberg (1976) (but see Roelfsema (2023); Shadlen & Movshon (1999) for alternative hypothesis), provides a comprehensive theory to understand how the brain integrates and perceives diverse sensory inputs. This theory asserts that synchronous neural activity, induced by neural oscillations, plays a foundational role in cognitive processes. According to this theory, when distinct features of a sensory stimulus are processed by specialized regions of the brain, the neurons responsible for representing these features synchronize their firing patterns at specific frequencies. This synchronization of neural oscillations enables precise coordination and the temporal binding of neuronal responses from different regions, thus uniting them into a coherent percept (Fries et al., 1997; 2002).
The concept of binding is also linked with the core principles of Gestalt psychology (Wertheimer, 1938), a field dedicated to understanding perceptual organization and how meaningful structures emerge from sensory data (Gray & Singer, 1989). Gestalt principles, such as proximity, similarity, and closure, highlight the brain’s innate tendency to organize sensory information into coherent and structured wholes rather than process isolated parts (Todorovic, 2008). Synchrony can therefore be viewed as a mechanism that induces the grouping and integration of sensory elements based on these Gestalt principles (Gray et al., 1989).
In summary, the binding by synchrony theory suggests that synchronized neural activity is crucial for binding distributed sensory responses. In this paper, we use complex-valued representations to mimic this synchrony mechanism and the Kuramoto dynamic to support Gestalt principles of proximity and similarity. In our model, we assume an oscillation at a single frequency (comparatively to the brain which can exhibit oscillations at different frequency bands) for simplicity purposes. The phase of the complex-valued neuron will, therefore, represent the phase with respect to an ongoing oscillation, and a group of neurons sharing the same phase value will be akin to a synchronized population.
4 KomplexNet: Kuramoto synchronized complex-valued network
In the following, we describe the details of KomplexNet’s implementation. We use a Kuramoto system at the first layer to synchronize the phases and apply operations in the complex domain for subsequent layers, and we finish with the addition of feedback connections affecting the phase synchrony.
4.1 Kuramoto dynamic
Kuramoto model.
Instead of considering synchrony as an emergent phenomenon, we choose to induce it using a Kuramoto system (Kuramoto, 1975) to organize the phases of the first layer. This process serves to synchronize the phases before incorporating them into the activity of the network. As we are dealing with complex-valued activity, we also need an amplitude value for each neuron at the first layer. This information is obtained by applying a real-valued convolution to the input image. To reach a synchronized state, we adapt the original equation of the Kuramoto model (Equation 3 in Kuramoto (1975)) and propose the dynamic described in Equation 1.
The core idea of the Kuramoto model is to synchronize a population of oscillators by mutual influence. In our case, the oscillators are the phases (with H and W denoting the size of the input image), where a phase will be influenced through the sine of the difference between itself and the rest of the population. This influence will be modulated by a learnable coupling kernel (with and ) together with a global desynchronization interaction , as well as the amplitude, , associated to each influencing phase. In other words, each phase synchronizes with its neighbors (defined by the spatial range of the kernel) and desynchronizes with phases further apart. At the population level, the system favors the emergence of several clusters of phases. Lastly, acts as a gain parameter, modulating the phase update at each timestep.
Learning the coupling kernel.
The coupling kernel is expected to capture the interactions and mutual influence between phases. A positive/negative value in this kernel means that the corresponding two neurons will tend to synchronize/desynchronize their phases. The kernel should learn the inherent structure of the objects in the dataset to adjust the interactions between nearby phases. For example, in the case of handwritten digits, the objects are mostly vertical, hence the kernel should favor positive interactions between phases along the vertical axis. Here, we initialize the kernel with a 2D Gaussian to encourage interactions with closer neighbors.
To learn this kernel, we use the cluster synchrony loss defined in Ricci et al. (2021):
where is the circular variance, the average of a phase group, and the number of groups. The first part of this loss measures the intra-cluster synchrony while the second part represents the inter-cluster desynchrony. Minimizing the loss resolves to minimize the variance inside groups (each phase cluster should have the same value) and minimize the proximity between the centroids of the clusters on the unitary circle (the clusters should cancel each other out).
In Figure 3, we show some visualization of the phases obtained after 15 steps of our Kuramoto model on the Multi-MNIST dataset (Sabour et al., 2017) (see also Fig. 9 to visualize the learned kernel). The plot represents the 8 convolution channels, masked by the intensity of the amplitude, with the color indicating the phase value. The phases from the same digits are synchronized, and the two clusters of phases are desynchronized with an almost opposite position on the unitary circle. Interestingly, the model does not learn to systematically affect one specific phase value to a class of digits (as observable in Figure 3 with two ’9’ represented by distinct colors/phase values in the first two images). Indeed, in the binding by synchrony theory, there is no requirement that a given object should always and systematically be assigned the same phase value, as this would seriously limit the flexibility and adaptability of the coding system.
4.2 Complex representations in an artificial neural network
Overall architecture.
Our model, KomplexNet combines the Kuramoto dynamic from Section 4.1 and additional complex operations described below, as represented in Figure 2 and Algorithm 1. To summarize, we start by extracting features from the input images by performing a real non-strided convolution (8 channels). The initial complex activity comprises these features along with random phases to perform one step of our Kuramoto dynamic. The resulting activation, , is propagated in a bottom-up manner through one strided complex-convolution () and two linear layers (respectively and ). This is repeated for several timesteps to allow the phases to reach a stable synchronized state through the Kuramoto dynamic.
Each step of the Kuramoto model outputs a new state of the phases, combined with the amplitude extracted via the first real convolution to instantiate the complex activity.
Complex operations.
We first redefine the standard operations of a convolutional network to be compatible with complex activations . We apply some biologically plausible transformations that allow the model to perform well, as used in (Reichert & Serre, 2013; Löwe et al., 2022; Stanić et al., 2023). We define the linear operations (convolutions and dense layers) as such:
The newly obtained activity results from applying real weights to both the real and imaginary parts of the input activity, modifying the amplitude and the phase jointly. Then, as first proposed by (Reichert & Serre, 2013), we apply the classic term, representing a gating mechanism that selectively weakens out-of-phase inputs, preventing inhibition (caused by a negative weight) from leading to the same result as desynchronization (phases in the opposite direction):
Lastly, we apply a ReLU non-linearity only on the amplitude. Considering that a complex amplitude is by definition positive, we start by normalizing it before applying the desired function, as done by (Löwe et al., 2022; Stanić et al., 2023). This last step ends a block of operations representing one layer in our model.
We can additionally observe, in Figure 4 the effect of the complex operations by visualizing the phase distribution at each layer of the model at the last timestep. The first row represents the polar distribution of the color-coded phases at each layer. We can see that the Kuramoto model allowed the phases to reach a state with two opposite clusters – one for each digit – and this distribution is conserved in all the subsequent layers. We show in the second row the same activity but detailing the spatial information provided by the convolutions of the first and second layers.
4.3 Implementing feedback
We finally propose an extension of KomplexNet by implementing feedback connections to influence the phase synchronization (dashed lines on the right in Figure 2). At each timestep, the higher-level representations from the latter layers (carrying information about advanced features, object parts, and object classes) can help the synchronization process of the first layer. We combine the lateral synchronization Kuramoto dynamic of with feedback synchronization from higher layers. The resulting phase update per timestep at the first layer is described analytically in Equation 6. The first line defines as the lateral synchrony in (same as Equation 1) and the second combines with the sum of the Kuramoto dynamics across layers.
We instantiate the activity of each layer at the first timestep through a feedforward propagation and introduce the feedback starting at the second step. Algorithm 2 details the corresponding update in the dynamic of the whole model. Similarly to the lateral coupling kernel of layer , the feedback coupling kernel coming from is defined in . However, the coupling matrices from the dense layers are defined in with representing the number of neurons in each layer. More specifically, the convolutional layer still provides spatially structured information to the phases of the first layer . Conversely, the phases of the dense layers affect all the phases of without spatial structure but provide generic information about the identity of the objects. The feedback couplings in the Kuramoto equation do not comprise the global desynchronization term (unwarranted since the latter layers are not spatially structured). However, we initialize all the feedback kernels around 0 (before training) to facilitate reaching negative coupling values, thus desynchronizing certain phases of with phases from higher layers when they do not encode for the same object.
4.4 Experimental setup
Datasets.
We perform our experiments on two different datasets. The first one is the Multi-MNIST dataset (Sabour et al., 2017), consisting of images containing two hand-written digits taken from the MNIST dataset. More specifically, we generate empty (black) images of size , downsample the MNIST images by a fixed factor (depending on the number of digits we want to fit in the image), and place the first digit at a random location around the upper-left corner. We then randomly pick a distinct second digit and assign it a position in the image under two constraints: not surpassing a predefined maximum amount of overlap, while still fully appearing in the image (none of the digits are cut by the image border). We generate a non-overlapping version of the dataset (maximum overlap = ) and an overlapping version where the digits can overlap up to of their active pixels. We use the same procedure to generate images containing more than two digits. Similarly, we generate a version of this dataset with greyscaled CIFAR10 (Krizhevsky et al., 2009) images in the background. This makes the digits less easy to separate for the Kuramoto model and represents a more ecological setting to test our models.
For both of these datasets, the models are evaluated on their ability to recognize and classify the two digits, out of 10 different possible classes. When generating the images, we also generate an associated mask tagging the different objects: first digit and second digit (the same logic potentially extends to a higher number of digits). When the digits overlap, the overlapping region is considered as an additional object. The background is not considered as an object. These masks are used to compute the cluster synchrony loss (Equation 2) and do not provide information about the identity (label) of the digits.
Baselines.
We compare both versions of our model (KomplexNet and KomplexNet with feedback) with different baselines to highlight our contributions: a real model (with an architecture and a number of parameters equivalent to KomplexNets) and a complex model without the Kuramoto synchrony (random phases at the first layer). We additionally show the performance of a complex model with an ideal phase separation as an upper baseline: for this model, using the ground-truth masks, we assign the phases by randomly sampling equidistant groups on the unitary circle and affecting each value to one object in the image (with the number of objects, not including the background). When digits overlap, we affect an intermediate phase value (circular mean of the two-digit values) to the overlapping pixels.
Model and training.
We train each family of models end-to-end using Adam (Kingma & Ba, 2014), a fixed learning rate of 1e-3, and a batch size of 128 or 32 depending on the dataset. KomplexNets are trained by accumulating the binary cross-entropy loss at each timestep and then combining it with the synchrony loss from the last timestep. The balance between the two quantities is modulated by a hyperparameter, as illustrated in Equation 7. All experiments are implemented in Pytorch 1.13 (Paszke et al., 2017) and run on a single NVIDIA V100. Each curve in the following plot represents the average and standard deviation over 50 runs with different random initializations. Complementarily, we present in the appendix the test accuracy of the best models on the validation set. To obtain the best hyper-parameter values, we run a hyper-parameter search and use the best combination of values out of 100 simulations. The concerned parameters are: the desynchronization term , the gain parameters for each layer , the coupling kernel sizes and , and the balance of the losses .
Evaluation.
As we are optimizing two different losses (cluster synchrony and classification), we systematically quantify the performance of our models compared to the baselines along those two separate axes. In all the following sections, we show the first loss under the synchrony label and use the performance label for the classification objective. The models are all trained with two or three non-overlapping digits. In the next sections, we present the results on in-distribution images (two-digit, non-overlapping images, all different from the training set) and then evaluate the robustness and generalization abilities of the models (without re-training or fine-tuning). We define here robustness as the model’s ability to perform the same task given an altered image (compared to the training distribution), while generalization represents the capability of the model to perform a slightly different task from the one it was trained on (here, the categorization of more or less digits than during training).
5 Results
5.1 In-distribution performance
Synchrony.
We start by evaluating the ability of the Kuramoto model to correctly separate the objects, compared to the other complex baselines (we cannot explicitly evaluate object separation for the real-valued baseline, as this requires phase information). In Figure 5 panel A, we quantify the quality of the solutions provided by the four models using Equation 2, with the random case as a lower baseline and the ideal case as an upper baseline; we see that the solutions obtained with Kuramoto (and especially with feedback connections) converge over time towards the ideal value. In panel C, we show a qualitative example of phase synchrony for an image taken from the test set (illustrated in panel A). As expected, when the digits are spatially separated on a clean image, the phases obtained with the Kuramoto model almost perfectly synchronize inside the digits and desynchronize between digits: the two clusters have opposite values on the unitary circle (polar plots in the bottom row) and all the pixels within one digit have similar phase values (phase maps across the 8 convolution channels, middle row). In this image, the area where the digits are close and almost touching is assigned a phase value in-between the two clusters; this seems a reasonable solution because the coupling kernel around this region will tend to synchronize the phases across the two digits. A smaller coupling kernel would avoid this problem, but it would then take more timesteps to reach a synchronized state. Visually, the obtained phase maps with KomplexNets seem to lie between the random-phase model (left sub-panel) and the ideal-phase model (right sub-panel) where the clusters show no variance even in ambiguous regions of the image. KomplexNet with feedback shows slightly less dispersed clusters than the feedforward KomplexNet model (middle two sub-panels).
Performance.
We then evaluate the effect of synchrony on the performance of the model. At each Kuramoto step, we evaluate the KomplexNets and obtain an accuracy curve over time. Conversely, the baselines are deprived of temporal dynamics and therefore yield only one accuracy value. The results are reported in Figure 5, panel B. We can observe that the KomplexNets’ performance starts below the real-model and random-phase baselines (because the phases are not synchronized yet) and out-perform them after 5 timesteps (for KomplexNet with feedback) or 7 timesteps (for KomplexNet without feedback). Surprisingly, after about 10 timesteps, KomplexNet reaches a performance on par with the ideal phase model, while KomplexNet with feedback even outperforms it. This observation sheds light on the phase-synchronization strategy we define as “ideal” here. Because the phase initialization takes place after a convolution, the resulting activity is spread over a slightly larger extent than the "ideal" mask (based on active image pixels). Because of this discrepancy, some activated neurons around the digits are not affected by the phase initialization process, introducing noise in the phase information that is later sent to the rest of the network. Conversely, the Kuramoto dynamic acts on all active neurons, potentially leading to more faithful phase information in the in-distribution case.
We provide in Appendix Figure 10 additional experiments testing the models on more timesteps than during training; the results indicate that the synchronized Kuramoto state is stable over time, and the associated performance improvements persist. We also test in Appendix Figure 12 the effect of the feedback coming from each layer (, , ) separately; every single layer provides an amelioration, but the model with feedback from all layers combined shows the greatest performance.
5.2 Robustness
We then evaluate the robustness of our trained models on out-of-distribution images. The task remains two-digit classification, and we evaluate both objectives (synchrony and performance) on images with overlapping digits or with additive Gaussian noise.
Synchrony.
Similarly to the previous section, we observe the cluster synchrony loss of KomplexNet with feedback reaching a lower (i.e. better) score compared to KomplexNet (see Figure 6, first row), for both the “overlap” and the “noise” conditions. In both cases but not surprisingly, the gap with the ideal phases remains higher than before, since the ground-truth phase masks (used for the “ideal-phase” model) are not affected by noise or by digit overlap (when digits overlap, pixels from the overlapping region are considered as a separate group in the ground-truth mask, but this group is not included in the synchrony loss computation). We provide a visualization of the phases for one image (the example shown in Figure 6) in the Appendix Figure 14.
Performance.
KomplexNet and KomplexNet with feedback show more robustness than the baselines, outperforming the real model in accuracy by 10 to 15% (Figure 6, second row). The real model and the random-phases complex model are very affected by the perturbations, while the ideal-phase complex model shows less altered performance. The accuracy values of KomplexNets lie between the upper baseline and the two other models and surpass them after fewer timesteps than in the previous case (Figure 5), showing how synchronized phases help resolve ambiguous cases. More specifically, KomplexNet with feedback remains more robust than KomplexNet, motivating the use of feedback connections to improve phase synchronization.
5.3 Generalization
To evaluate our models’ generalization abilities, we report in this section their synchrony and performance when trained on either two or three digits in the images and then tested on the same or a different number of digits, from two to nine digits.
Synchrony
Figure 7 illustrates the generalization abilities of the KomplexNets at synchronizing the phases in this out-of-distribution setting. Interestingly, despite not having seen 3 digits during training, the coupling kernel of the model trained on two digits can create a third cluster, equidistant to the others on the unitary circle, and correctly affect a single phase value per digit. Likewise, the model trained on three digits adapts well to the two-digit case: the model doesn’t create a third phase cluster but only shows higher variance in the two opposite clusters compared to the model trained on two digits.
The evolution of the cluster-synchrony losses illustrates well the general case: losses for both KomplexNets start around the random-phase case and end close to the ideal-phase case. Interestingly, for a given test setting, no matter the training setting (same or different digit number), the models reach approximately the same synchrony values at test time. This observation highlights an additional form of robustness of our models and suggests an object representation ability not present in non-complex and non-synchronized (random-phase) complex models.
Performance
In the same way, we report the classification accuracy of the models and their baselines when trained on two or three digits and tested on two or three digits (Figure 8, panel A). Consistent with the previous results, both KomplexNets outperform the real and random-phases model baselines, both when testing in-distribution and out-of-distribution. More interestingly, both versions of KomplexNet reach almost the same performance on each given test set (within ), no matter the number of digits seen during training. Conversely, the baselines (real and random-phase models) suffer more from this change, leading to a very consequent gap in performance (up to ) on the off-diagonal (out-of-distribution testing) plots.
Given this success in generalizing the classification task to one more or one less digit compared to the training set, we next evaluate the maximum number of objects that can be handled by KomplexNets. Consequently, we report the performance of all the models on two to nine digits in the image. In Appendix Figure 13, we report the absolute performance of the models at the last timesteps for each test set (one test set corresponding to a fixed number of digits in the images). As could be expected, performance decreases rapidly as the number of simultaneous objects to classify grows from two or three digits (the number that the models were trained on) up to 9 digits. However, performance drops more quickly for the baselines (real and random-phase models) than for KomplexNets. As the exact advantage of KomplexNets is hard to quantify from this plot, we also report in Figure 8 the difference in performance between all the models and KomplexNet (panel B), and all the models and KomplexNet with feedback (panel C), when trained on either two (first column) or three digits (second column). On these plots, zero difference means that the tested model reaches the same performance as KomplexNet (for panel B, or KomplexNet with feedback for panel C), while a negative difference means that the tested model performed worse than KomplexNet (and conversely). In panel B, we observe a negative difference between KomplexNet and the baselines (except KomplexNet with feedback), persisting up to 8 digits with a peak around 3 to 5 digits. Similarly, KomplexNet with feedback (panel C) systematically outperforms all the other models from two to 8 digits. The nine-digit test set is very hard for all the models because it is the furthest one from the in-distribution case and, as observable in Figure 13, accuracy is very low.
Overall, these results reveal that phase synchronization makes the models more robust and general by rendering them less sensitive to out-of-distribution shifts.
5.4 Additional experiments
Experiments on non-uniform background.
To demonstrate that our method is not restricted to images with empty backgrounds, we train our models and the baselines on a version of the multi-MNIST dataset with randomly drawn CIFAR images in the background. The results are presented in Figure 15. Because the meaningful information contained in the image is more difficult to extract, both synchrony and performance measures are affected for all the models. However, KomplexNet and especially KomplexNet with feedback still clearly outperform the baselines. This additional experiment highlights the robustness of the Kuramoto model: the difficulty is in ignoring noisy information in the background, therefore relying less on proximity and more on similarity principles. For this reason, the cluster synchrony loss of KomplexNet is very altered and far from the ideal case, but the feedback connections of KomplexNet with feedback help to bridge the gap.
Leveraging the temporal dynamic from temporal inputs.
Finally, we evaluate the advantage of the Kuramoto dynamic on dynamic inputs. More specifically, we investigate whether the temporal dimension of the Kuramoto dynamic can act as a memory mechanism when the input is transformed from static images to videos (with digits moving across the frames). We show in Figure 16 the accuracy per timestep relative to the frame with the maximum amount of overlap (denoted timestep 0). Compared to the real baseline, as well as KomplexNets tested on each frame separately, we observe a significant increase in accuracy when the models are tested on the moving object test set. These results suggest that the phase information coming from the previous frames (where the digits overlapped less) is maintained, leading to better performance. Overall, it confirms the hypothesis that Kuramoto can act as a system with memory, able to use information from the previous timesteps to create a more qualitative phase separation than a system deprived of such context.
6 Conclusion
6.1 Summary
Here, we propose a model combining complex-valued activations with a Kuramoto phase-synchronization dynamic modeling of the binding-by-synchrony theory in neuroscience. A complex-valued neuron can simultaneously indicate the presence of a feature by its activation amplitude (just as in standard real-valued neural networks), and tag the group or object to which this feature belongs by its activation phase. The Kuramoto model serves as a synchronization process, where the coupling kernels implement Gestalt principles of proximity and similarity and act as an inductive bias. We show that our model outperforms its real counterpart as well as a complex-valued model with random phases on multi-object recognition. More interestingly, the phases play an important role in introducing some notions of object representation in the models, making them more robust to ambiguous cases such as overlapping digits, noisy images, more digits in the images, etc.
We additionally propose a way to introduce feedback connections in the model, acting on the phases by using information from higher layers to enhance synchrony. As the phase information becomes more reliable (better cluster synchrony score) with the feedback extension, the performance, robustness, and generalization of the model increase, highlighting the added value of good phase synchronization for multi-digit classification.
6.2 Limitations and discussion
Model Depth and Scalability
The models studied in this work are relatively shallow. We intentionally designed a limited architecture to restrict feature representation abilities and highlight the benefits of synchrony in such a setting. While this serves as a proof of concept, scaling up our approach requires identifying tasks and datasets where state-of-the-art models struggle with feature binding in multi-object scenarios. Given the computational demands of training large models on extensive datasets, we opted for a smaller-scale demonstration.
Hyper-Parameter Dependence
The Kuramoto dynamics introduced in our model rely on several hyper-parameters, including , for each layer , the coupling kernel sizes and , and the loss balance parameter . Finding optimal values for these parameters requires hyper-parameter tuning to ensure phase synchronization. However, our experiments indicate that these values remain consistent across different tasks (see Table 1), suggesting that hyper-parameter tuning is only necessary when changing datasets.
Explicit vs. Emergent Synchronization
A potential limitation of our approach is the explicit addition of Kuramoto dynamics, in contrast to prior work where synchrony emerges through training (Löwe et al., 2022; Stanić et al., 2023). As noted by Stanić et al. (2023), the Complex Autoencoder (CAE) (Löwe et al., 2022) can achieve phase synchronization, but only on simple datasets with a limited number of objects. While Stanić et al. (2023) proposed a method to scale this behavior, they required an additional objective to achieve synchronization on more complex datasets. The emergence of synchrony at scale for complex visual scenes remains an open and non-trivial challenge. Our approach circumvents this difficulty by directly incorporating synchrony via Kuramoto dynamics, providing a controlled framework to assess its benefits. This serves as a proof of concept, demonstrating the advantages of complex-valued representations and synchronized activity in visual tasks. Furthermore, our approach aligns with experimental findings in neuroscience (Fries et al., 1997), emphasizing the role of synchrony in early visual processing.
Neuroscientific Relevance
Our study is primarily motivated by the binding by synchrony theory, a concept widely discussed in neuroscience. While this theory remains influential, it has faced experimental challenges (Roelfsema, 2023; Shadlen & Movshon, 1999). However, recent evidence supports the functional importance of synchronized activity (Fries, 2023). Our work does not aim to provide new insights into biological vision but instead proposes synchrony as a viable solution for feature binding in artificial models. By leveraging neuroscientific principles, we offer an alternative computational mechanism that may inspire future advancements in deep learning architectures.
7 Acknowledgment
Our work is supported by ONR (N00014-24-1-2026), NSF (IIS-2402875) to T.S. and ERC (ERC Advanced GLOW No. 101096017) to R.V., as well as “OSCI-DEEP” [Joint Collaborative Research in Computational NeuroScience (CRCNS) Agence Nationale Recherche-National Science Fondation (ANR-NSF) Grant to R.V. (ANR-19-NEUC-0004) and T.S. (IIS-1912280)], and the ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A-0004) to R.V. and T.S. Additional support was provided by the Carney Institute for Brain Science and the Center for Computation and Visualization (CCV). We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program as well as computing hardware supported by NIH Office of the Director grant S10OD025181.
Appendix A Appendix
A.1 Algorithms
We detail here the different steps of both versions of KomplexNet using the pseudo-code algorithm. Algorithm 1 describes the operations of KomplexNet and Algorithm 2 specifies how feedback connections integrate in the previous dynamic.
A.2 Coupling kernel
A.3 Complementary results
We show in this section the complementary results and various tests, to give more insights into the functioning of the proposed method.
Test on more timesteps.
Without retraining KomplexNets, we test the models on more timesteps than during training. The resulting (Figure 10) performance shows the robustness of the Kuramoto dynamic to maintain a good phase representation over time.
Best validation model.
The results presented in the main paper are obtained by averaging the performance (synchrony and accuracy) on 50 different initializations and completing with the standard deviation. Complementary to this choice, we select the best model on the validation set and plot in Figure 11 the synchrony on the first row and the performance on the second row for each type of model and three different test sets (in-distribution, overlap, and noise). The results are consistent with the ones in the main text (except for the performance of KomplexNet with feedback on noisy images).
Influence of the feedback layer.
KomplexNet with feedback receives feedback information from all the layers. To highlight the contribution of each layer specifically, we train separated models with feedback coming from only one of the layers and we report the performance in Figure 12. We can first observe that KomplexNet with feedback (coming from all the layers) outperforms all the other versions on robustness tests, emphasizing the need for various types of information to solve ambiguous cases. Conversely, KomplexNet performs worse than all the models with feedback (both in-distribution and robustness). We additionally provide for each case the value of the hyper-parameters found to maximize the performance in Table 1. In conformance with our expectations, has to be higher when the models are provided with feedback connections to compensate for more synchronized activity coming from other layers. The size of the local kernel remains the same across the versions, corresponding more or less to the size of the digit in the image. Likewise, is smaller than to account for the downsampling of . Finally, the influence of local synchrony, modulated by is higher than the influence of the feedback phases ( for ).
| KomplexNet | 0.2 | 13 | - | 0.006 | - | - | - |
| KomplexNet with feedback from l2 | 0.4 | 13 | 5 | 0.006 | 0.003 | - | - |
| KomplexNet with feedback from l3 | 0.5 | 13 | - | 0.005 | - | 0.004 | - |
| KomplexNet with feedback from l4 | 0.5 | 13 | - | 0.005 | - | - | 0.004 |
| KomplexNet with feedback from l2,3,4 | 0.5 | 13 | 5 | 0.009 | 0.005 | 0.004 | 0.004 |
Raw performance on the generalization test sets.
The main text reports the difference in the performance of KomplexNet and KomplexNet with feedback when tested on two to nine digits to evaluate their generalization abilities. We show in Figure 13 the raw performance of each model and the baselines trained on two or three digits. All the models show more difficulty in classifying correctly as the number of digits in the image increases. However, KomplexNets show consistently a higher performance (see Figure 8).
A.4 Visualizations
Similarly to the in-distribution tests, we provide in Figure 14 visualizations of the phases of all complex models. The random case remains unchanged. However, we can observe that KomplexNets find a solution for the overlapping digits by almost creating three equidistant clusters, with the overlapping pixels belonging to a cluster in between the two others on the unitary circle. Despite no supervision, KomplexNets found a solution close to the one we adopted for the “‘ideal” case, consisting of affecting the overlapping pixels to a third cluster, but keeping the two other clusters opposite to each other. Additionally, when we add Gaussian noise to the images, the models struggle more to separate the digits, but KomplexNet with feedback remains able to affect opposite values to the digits (the clusters just show more variance).
Phases obtained on robustness tests.
A.5 Additional experiments
We finally show the results corresponding to the additional experiments detailed in the main text with variants in the datasets to show different uses of KomplexNet.
CIFAR10 images in the background.
We create an additional dataset with the same digits but with the background filled with RGB content, namely images from the CIFAR10 dataset. We retrain the models on the new images (the task remains unchanged: 2-digit classification) and report the cluster synchrony loss and accuracy on the test set in Figure 15. Compared to the previous datasets (with a uniform background), the models achieve a lower performance. Indeed, the task is now harder due to less salient information to extract. This version is particularly harder for the Kuramoto model: the propagation of phase synchrony is enhanced by the activated background, even though the information is irrelevant. For this reason, the cluster synchrony is much higher than before and very far from the ideal scenario. As a result, the accuracy is also further from the ideal case. However, KomplexNets are still better than the baselines, confirming that the Kuramoto dynamic remains helpful on RGB images. More specifically, the feedback connections show a clear advantage on the cluster synchrony loss compared to KomplexNet without feedback, resulting in a slight increase in test accuracy.
Moving digits.
The second variant in the dataset is to convert static images into videos with both digits moving across frames. With this version, we aim to evaluate whether we can use the Kuramoto dynamic along with the phase information as a memory mechanism to help resolve ambiguous cases (overlap between digits in this case). To test such a hypothesis, we generate videos of 30 frames, starting from the two digits overlapping (with a maximum of 25% of the active pixels) in the center of the image. We then generate a random trajectory for the first digit and use the opposite trajectory for the second. We let the objects move along these trajectories for 15 frames (making them bounce on the border of the image to prevent them from disappearing and controlling for the amount of overlap at each frame) and use these frames as the first half of the video. We take the opposite trajectories and apply the same procedure to generate the 15 other half of the video. In other words, the resulting videos start with both digits at a random location, then slowly move closer to each other, reaching a maximum amount of overlap in the middle before moving away from each other. We are interested in the performance of the model around this frame of maximum overlap (10 frames before and after). We use the first frames to let the Kuramoto model converge to a clean phase separation. We evaluate KomplexNet with and without feedback on the videos, as well as the same models tested on each frame separately. For the models tested on the videos, the task has the additional difficulty of adapting the phase information to moving information: each frame corresponds to a single Kuramoto step, potentially leading to noisy artifacts in the phase information because the models were trained on static images. However, the models tested on each frame separately, despite having the time to converge on clean phase separation, do not have access to frames with less overlap between digits. We show in Figure 16 (left panel) the raw performance 10 frames before and after reaching the maximum overlap (denoted frame 0) for KomplexNet with (purple) and without feedback (red) given the two types of input (static in dashed lines and videos in plain lines). We report the performance of the real model (tested on static images only because of the absence of temporal dynamic in the model) to confirm the benefit of KomplexNet on this version as well. We present the videos from Frame 0 to Frame 30 as well as from Frame 30 to Frame 0 to correct for potential bias in the generation of the dataset. On the left Panel, we observe that the models which were presented the moving digits are outperformed by the models tested on single images at the first timesteps. This is explained by the fact that the digits do not overlap (or barely do) 10 frames before Frame 0. Therefore, given several timesteps to converge on the frame, the "static" models reach a cleaner phase separation leading to better accuracy. However, as we get closer to the maximum overlap, the "static" models show a drop in performance, compared to the "dynamic" models suggesting the use of the phase separation coming from the frames where the digits did not overlap. Finally, as the digits get away from each other, the "static" models recover their performance, outperforming after a few frames the dynamic model. This result is shown differently on the center panel. We show here the difference in the accuracy of each model from Frame 0 to Frame 30 with the accuracy from Frame 30 to Frame 0. Every "static" model has a null difference because the dataset was symmetrized. However, the "dynamic" models show a positive difference before Frame 0 and a negative difference after, suggesting that they use the clean phase from the previous steps to keep a good phase separation when the digits overlap. However, they take a few steps to recover from the phase corruption induced by a lot of overlap between digits. Finally, we show on the right panel the difference in performance between static and dynamic models. We can observe a maximum difference at Frame 0 in favor of dynamic models, outperformed by their static counterparts when the amount of overlap is reduced. These results confirm the possible use of the Kuramoto dynamic as a mechanism of memory, helping resolve extremely overlapping cases and being compatible with dynamical inputs.